OpenStudy (hpfan101):

Differentiate: \[f(x)=\arctan(\frac{ x-1 }{ x+1 }) \]

OpenStudy (hpfan101):

\[f'(x)=\frac{ 1 }{ 1+(\frac{ x-1 }{ x+1 })^2 }\times(\frac{ x-1 }{ x+1 })'\]

OpenStudy (hpfan101):

I'm not sure how to simplify the expression after taking the entire derivative.

OpenStudy (freckles):

for the fraction there you could multiply top and bottom by (x+1)^2 giving you... \[f'(x)=\frac{(x+1)^2}{(x+1)^2+(x-1)^2} \cdot (\frac{x-1}{x+1})'\]

OpenStudy (hpfan101):

Oh, ok, thanks! I think I know what to do for the rest of the problem.

OpenStudy (freckles):

your answer should turn into a really pretty answer at then end...

OpenStudy (freckles):

Here is a hint: it will be of the form: \[\frac{a}{bx^2+c} \\ \text{ where } a,b,c \text{ are constants }\] if you do not end up with this form please show me what you have

OpenStudy (hpfan101):

Ok, will do.

OpenStudy (hpfan101):

My final answer is: \[f'(x)=\frac{ 1 }{ x^2+1 }\]

OpenStudy (freckles):

yep looks great!

OpenStudy (freckles):

a=b=c=1

OpenStudy (hpfan101):

Cool, thanks for the help!